OpenStudy (anonymous):

The first five terms in a quadratic sequence are shown below. 5, 14, 29, 50, 77,... What is the nth term in the sequence? A. 4n2 + 2 B. 3n2 + 3 C. 3n2 + 2 D. 2n2 + 3

OpenStudy (anonymous):

So take the first term i.e 5, and the second term is 14

OpenStudy (anonymous):

for first term n=1 and for second term n=2

OpenStudy (anonymous):

as you see the options are in the for an^2 + c where the values of a and c are changing

OpenStudy (anonymous):

so for n = 1, the quadratic will be a*1^2 + c = a+c for n=2, it will be a*2^2 + c = 4a+c

OpenStudy (anonymous):

we know a+c = 5 (1st term) and 4a+c = 14 (2nd term)

OpenStudy (anonymous):

So if we subtract the first term from 2nd term then we get (4a+c) - (a+c) = 14 - 5 So 3a = 9

OpenStudy (anonymous):

Now you can find the value of a once you get use the value of "a" and substitute in the equation a+c = 5 to get the value of c

OpenStudy (anonymous):

Now see which option satisfies

OpenStudy (anonymous):

It would be B wouldn't it..

OpenStudy (anonymous):

Nope but you are partially correct,

OpenStudy (anonymous):

the 3n^2 part is correct. Did you get what I said earlier?

OpenStudy (anonymous):

I honestly understand now. Thank you so much for your help, now that you explained it better then my teacher, I now understand this.

OpenStudy (anonymous):

Haha thanks,so what is the answer?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

Haha Thanks a billion!